### v2

@@ 3.4 Implementation in MC/DC
MC/DC [12] is designed to implement and test new time-dependent Monte Carlo numerical methods at scale [40]. It uses a novel (for the field) development structure in which Python compute kernels are compiled via the Numba just-in-time compiler to run on CPUs and with the Harmonize runtime manager [41, 42, 43] to run on GPUs. MC/DC enables rapid experimentation with these methods at scale on both CPUs and GPUs with time-dependent problems of interest.
@@
This process is similar to that used to implement hybrid surface-delta tracking methods in MCATK [8]. However, previous work in MCATK implemented a hybrid delta tracking algorithm, which was conceived in part to integrate into MCATK smoothly.
@@ 4 Verification of delta tracking with continuously moving surfaces
To verify that delta tracking can be used in conjunction with continuously moving surfaces in MC/DC, we use the moving pellet test from MC/DC’s regression test suite [12]: a cylindrical fuel element moves through a region that also has a small source. As the pellet moves closer to, then further away from, the source region, the fission rate in the pellet changes. Figure 8 at top shows the fission reaction rate density at various points in time. The outline of the pellet is clearly visible in a number of time steps. These plots were produced using delta tracking with a collision estimator (to compute fission reaction rates).
@@ 5 Benchmark Problems
revOneWe explore these methods using time-dependent fixed-source problems. “Fixed source” and “external source” refer to a class of problems that are not eigenvalue problems. In fixed source problems, the solution has a magnitude that depends on the prescribed source. Neutron multiplication is allowed, provided the system is subcritical, and fixed source problems may or may not be time-dependent. Eigenvalue ($k$ or $\alpha$) problems only involve sources that depend on the solution. This work does not discuss how the implemented hybrid delta tracking methods impact $k$-eigenvalue or $\alpha$-eigenvalue calculation performance.
@@
revOneModified Dragon Burst $4.0\text{\times}{10}^{6}$ $3\text{\times}{10}^{9}$ CE (3 materials)
revOneModified C5G7 $3.9\text{\times}{10}^{6}$ $1\text{\times}{10}^{7}$ MG (7 groups, 8 delayed neutron precursor)
@@ 5.1 Kobayashi problem
We first consider a time-dependent version of the Kobayashi problem [44] introduced by Variansyah et al. [40] (a full description is provided by Variansyah [45]). We use 10 batches with ${10}^{9}$ particles per batch (${10}^{10}$ particles total) with surface tracking, standard delta tracking with a collision estimator, and a delta tracking using the voxelized tally method. This problem contains two materials, a low mean-free-path region, characterizing a solid ($\Sigma=$$0.1\text{\,}{\mathrm{cm}}^{-1}$), and a high mean-free-path region modeled by air, characterizing a void ($\Sigma=$$1\text{\times}{10}^{-4}\text{\,}{\mathrm{cm}}^{-1}$), in a single energy group. There are $1.20\text{\times}{10}^{5}$ structured tally voxels in the $x$-$y$ plane and in time. For this problem, we expect the voxelized tally method to perform better than standard delta tracking and surface tracking.
@@
Figure 9: Schematics for benchmark problems: (left) C5G7 reactor quarter (via reflecting boundaries) [39], (right) Dragon burst problem [46].
5.2 revOneModified C5G7 problem
@@
Figure 9 on the left shows the geometry for the next two simulations, which are based on the C5G7 benchmark problem [39]. In energy we use seven energy groups and eight delayed neutron precursor groups. We model a four-phase accident in which a pressurized light water reactor increases in power via the removal of control rods revOne[47]. The fixed-source source in the problem is located in throughout the active center pin of assembly four with a zero initial condition. MC/DC does not include any treatments for cross section heating during a simulation so this kind of multi-physical calculation not considered in this work. Figure 10 on the top shows the normalized flux density as a function of time (produced from point reactor kinetics) and the four phases shaded gray, green, red, and blue, respectively. Phase one (shaded gray) starts with the reactor operating in steady state. In phase two (shaded green), the control rods are removed from the reactor to cause the power to increase to a new steady state. Phase three (shaded red) begins when a bank of control rods gets stuck in the fully withdrawn position. Toward the end of phase three, the reactor sees a rapid increase in power that ends at $15\text{\,}\mathrm{s}$ when all control rods are forced into the reactor, ending the accident. In the fourth and final phase, reactor power decays as the delayed neutron population dies out. Figure 10 on the bottom shows plots of average flux in the $x$-$y$ plane at various points in time, including at $14.95\text{\,}\mathrm{s}$, the maximum power excursion. These results are from a simulation with ${10}^{6}$ particles in $10$ batches (a total of ${10}^{7}$ particles).
@@
Figure 10: revOneModified C5G7 stuck rod accident simulation: (top) flux densities through time showing the four phases shaded as gray, green, red, and blue respectively; (bottom) scalar flux on $x$-$y$ plane (top view) at points in time.
@@
We first modeled this problem using the seven-group materials in the C5G7 benchmark description [39], which we call C5G7 in the remainder of this work. We use $3.9\text{\times}{10}^{6}$ mesh cells in a 3D, time- and energy-dependent tally mesh. The movement of control rods into and out of the reactor is modeled with the continuous movement functionality of MC/DC [48]. To verify the delta tracking methods we explore convergence to the correct solutions using continuously moving surfaces, and we compare delta tracking solutions to solutions provided by surface tracking. This further verifies that delta tracking can be used in conjunction with continuously moving surfaces. Performance data is collected using ${10}^{6}$ particles in $10$ batches (total of ${10}^{7}$ particles).
@@ 5.3 C5CE problem
Next, we define a continuous-energy version of the previous problem (named C5CE), in which the reactor undergoes the same four-phase accident. revOneWe use the same C5G7 geometry [47], source distribution, and initial condition but establish each material using continuous energy cross sections. Table LABEL:tab:c5ce in Appendix B includes the material compositions for the C5CE problem. Using this benchmark, we evaluate both the voxelized tally method and the hybrid-in-energy method described in Section 3.3. Figure 1 shows the macroscopic total and majorant cross sections as a function of energy for the materials in C5CE. The neutron resolved-resonance region ends around $10\text{\,}\mathrm{keV}$, so we set the transition energy at $50\text{\,}\mathrm{keV}$. Particles moving at energies above $50\text{\,}\mathrm{keV}$ will use delta tracking, and those below will use surface tracking.
@@
5.4 revOneModified Dragon-burst problem
Conducted in 1944 during the Manhattan Project, the Dragon-burst experiments [46] proved that criticality and supercriticality could be achieved with prompt neutrons alone. Previous experiments, such as the Chicago Pile One, depended on the presence of delayed neutrons to achieve criticality. Figure 9 on the revTworight shows a schematic of the experiment where a highly enriched (75%) uranium hydride slug (UH10) was dropped through a tamper with additional fuel. The slug moves through the core, triggering a prompt supercritical reaction. The reactivity excursion ends when gravity causes the slug to fall out of the core. Kimpland et al. showed that this burst criticality experiment achieved an increase in the neutron population of more than nine orders of magnitude [46]. revOneThe total time of the reaction is fast enough such that cross section heating can be neglected [46]. Here we consider a less-reactive version of the Dragon burst problem (25% enrichment) to test delta tracking with the hybrid-in-material method described in Section 3.2, and to provide additional verification for delta tracking with continuously moving surfaces. Figure 11 shows the overall flux density as a function of time on the top and a $y$-$z$ plot of the scalar flux at various points in time on the bottom.
Figure 11: revOneModified Dragon burst simulation: (top) flux density as a function of time, (bottom) scalar flux on $y$-$z$ plane (side view) at points in time.
@@
We expect delta tracking algorithms to perform quite poorly in this model for a number of reasons. First, when undergoing delta tracking, particles in the air region will experience a large number of rejected virtual collisions, since the majorant is orders of magnitude larger than the cross section of air. Second, as much of the problem involves a near-void material, the collision estimator typically required by delta tracking will generate poor tallies in those regions. Third, the problem is geometrically simple, consisting of a rectangular slug moving through a rectangular slab with a properly sized hole such that the slug can move through the slab. The benefit of delta tracking over surface tracking is most pronounced in problems with many surface crossings [3]. Our objectives of comparing delta tracking methods for this specific problem are to confirm that delta tracking works with such a dynamic problem with continuously moving materials and to assess whether the hybrid-in-material method outperforms standard delta tracking.
@@ 6 Results
In this section, we discuss the performance results of the benchmarks modeled with surface, delta, and voxelized delta tracking, as well as the hybrid-in-energy and hybrid-in-material tracking algorithms. We also verify that the various delta tracking methods converge to the same solution as surface tracking when transporting on a geometry with continuously moving surfaces. All data, plots, the exact version of MC/DC, and input decks are available from Zenodo [1].
@@ 6.1 Performance results
All results presented below are executed on the Dane and Lassen high-performance computing systems at Lawrence Livermore National Laboratory (LLNL). Dane is a CPU-only system with dual-socket Intel Xeon Sapphire Rapids CPUs, each with 56 cores for a total of 112 CPU cores per node. Lassen is the open collaboration sibling to the Sierra machine with four Nvidia Tesla V100 GPUs and two IBM Power 9 CPUs per node. We make all performance statements with respect to a whole node of Dane (112 MPI threads, CPU) and a whole node of Lassen (4 MPI threads, GPU). We compile to CPUs with Numba v0.60.0, and compile to Nvidia GPUs with CUDA v11.8 and Nvidia-PTX with Numba v0.59.0666Numba v0.59.1 is the most recent version to support Power9 CPUs.. The delta tracking methods are built using MC/DC v0.12.0 [34], compiling on GPUs with Harmonize v0.0.2 [49]. All floating-point operations use double precision.
@@
The performance of the methods on the revOnemodified Dragon problem is at odds with observations of the algorithms on the other problems. Delta tracking simulations (both standard and voxelized) do not reach the end of the specified end time on Dane or Lassen in the eight hours allocated, on either machine. On Dane, as predicted, the hybrid-in-material method improves performance over the other two methods, as it completes the simulation in about four hours. However, standard surface tracking requires only one hour for this simulation. The hybrid-in-material method did not finish on Lassen. These observations are consistent with the understanding that delta tracking is typically less effective for problems with fewer surfaces and significant material heterogeneities.
@@ 7 Discussion, Conclusions, and Future Work
We have also demonstrated a novel hybrid-in-energy surface-delta tracking scheme, where surface and delta tracking are used at low and high energies, respectively. On the continuous energy version of the C5G7 benchmark geometry undergoing a four-phase transient, we observe an order of magnitude increase in FOM with the hybrid-in-energy scheme, on both CPU and GPU nodes. We have also confirmed that hybrid surface-delta tracking methods can yield improved results in problems with significant void regions [10].
@@
The main takeaway of this work is that delta and surface tracking do not have to be treated as discrete choices in the numerical method, as corroborated by papers from the developers of Serpent2 and MONK/MCBEND  [5, 11]. Greater performance can potentially be achieved by mixing and matching the underlying transport methods, given the physical parameters of a simulation and the relative strengths and weaknesses of a given transport application. As discussed in Section 3.4, implementing delta tracking in a surface tracking code is relatively simple, given a method to calculate a macroscopic majorant cross section.
@@
revOneMC/DC does not currently treat the unresolved resonance region. Work is ongoing to support unresolved resonances; when implemented the energy cutoff for the hybrid-in-energy delta tracking method could use the average value of the cross section to determine a cutoff value. For the majorant determination in the unresolved resonance region, the maximum cross section value in the unresolved resonance table will be used. This work only includes analysis for fixed source time-dependent problems. We leave the analysis of $k$ or $\alpha$ eigenvalue calculations for future work.
Methods for reducing the dimensionality of the majorant are also being explored to achieve a more-efficient majorant cross section lookup in distance to collision operations. Experiments with the collision estimator for scalar flux and methods of tallying into vacuum and low interaction rate regions are being considered, including implementing a cutoff method [10] or weighted delta tracking [4].
@@ Appendix A Continuous Energy Macroscopic Majorant
To compute a unified energy grid per material, we combine the energy grids from all nuclides in a given material. Then we call numpy.unique(), which returns a sorted array (from smallest to largest) with no repeating elements [50]. To compute a unified energy grid for the whole problem, we do the same but with all the nuclides in the entire problem.
Computing a macroscopic majorant for nuclides that are not on a unified energy grid requires two levels of interpolation to put a given macroscopic total cross section on a unified energy grid. First, interpolae from each nuclide’s microscopic cross section onto a material’s unified energy grid to compute a macroscopic total cross section. Then, perform a second interpolation from the material to the majorant’s unified energy grid. We use scipy.interpolate.interp1d() to interpolate from one energy grid to the next [51].

### v3

@@ 3.4 Implementation in MC/DC
MC/DC [morgan_monte_2024] is designed to implement and test new time-dependent Monte Carlo numerical methods at scale [variansyah_mc23_mcdc]. It uses a novel (for the field) development structure in which Python compute kernels are compiled via the Numba just-in-time compiler to run on CPUs and with the Harmonize runtime manager [brax2023, morgan_2025_monte, cuneo_2025_harmonize] to run on GPUs. MC/DC enables rapid experimentation with these methods at scale on both CPUs and GPUs with time-dependent problems of interest.
@@
This process is similar to that used to implement hybrid surface-delta tracking methods in MCATK [morgan_2021_mcatk]. However, previous work in MCATK implemented a hybrid delta tracking algorithm, which was conceived in part to integrate into MCATK smoothly.
@@ 4 Verification of delta tracking with continuously moving surfaces
To verify that delta tracking can be used in conjunction with continuously moving surfaces in MC/DC, we use the moving pellet test from MC/DC’s regression test suite [morgan_monte_2024]: a cylindrical fuel element moves through a region that also has a small source. As the pellet moves closer to, then further away from, the source region, the fission rate in the pellet changes. Figure 8 at top shows the fission reaction rate density at various points in time. The outline of the pellet is clearly visible in a number of time steps. These plots were produced using delta tracking with a collision estimator (to compute fission reaction rates).
@@ 5 Benchmark Problems
We explore these methods using time-dependent fixed-source problems. “Fixed source” and “external source” refer to a class of problems that are not eigenvalue problems. In fixed source problems, the solution has a magnitude that depends on the prescribed source. Neutron multiplication is allowed, provided the system is subcritical, and fixed source problems may or may not be time-dependent. Eigenvalue ($k$ or $\alpha$) problems only involve sources that depend on the solution. This work does not discuss how the implemented hybrid delta tracking methods impact $k$-eigenvalue or $\alpha$-eigenvalue calculation performance.
@@
Modified Dragon Burst $4.0\text{\times}{10}^{6}$ $3\text{\times}{10}^{9}$ CE (3 materials)
Modified C5G7 $3.9\text{\times}{10}^{6}$ $1\text{\times}{10}^{7}$ MG (7 groups, 8 delayed neutron precursor)
@@ 5.1 Kobayashi problem
We first consider a time-dependent version of the Kobayashi problem [Kobayashi2001] introduced by Variansyah et al. [variansyah_mc23_mcdc] (a full description is provided by Variansyah [variansyah_2025_15069882]). We use 10 batches with ${10}^{9}$ particles per batch (${10}^{10}$ particles total) with surface tracking, standard delta tracking with a collision estimator, and a delta tracking using the voxelized tally method. This problem contains two materials, a low mean-free-path region, characterizing a solid ($\Sigma=$$0.1\text{\,}{\mathrm{cm}}^{-1}$), and a high mean-free-path region modeled by air, characterizing a void ($\Sigma=$$1\text{\times}{10}^{-4}\text{\,}{\mathrm{cm}}^{-1}$), in a single energy group. There are $1.20\text{\times}{10}^{5}$ structured tally voxels in the $x$-$y$ plane and in time. For this problem, we expect the voxelized tally method to perform better than standard delta tracking and surface tracking.
@@
Figure 9: Schematics for benchmark problems: (left) C5G7 reactor quarter (via reflecting boundaries) [hou2017oecd], (right) Dragon burst problem [kimpland_2021_dragon].
5.2 Modified C5G7 problem
@@
Figure 9 on the left shows the geometry for the next two simulations, which are based on the C5G7 benchmark problem [hou2017oecd]. In energy we use seven energy groups and eight delayed neutron precursor groups. We model a four-phase accident in which a pressurized light water reactor increases in power via the removal of control rods [c5g7_zenodo]. The fixed-source source in the problem is located in throughout the active center pin of assembly four with a zero initial condition. MC/DC does not include any treatments for cross section heating during a simulation so this kind of multi-physical calculation not considered in this work. Figure 10 on the top shows the normalized flux density as a function of time (produced from point reactor kinetics) and the four phases shaded gray, green, red, and blue, respectively. Phase one (shaded gray) starts with the reactor operating in steady state. In phase two (shaded green), the control rods are removed from the reactor to cause the power to increase to a new steady state. Phase three (shaded red) begins when a bank of control rods gets stuck in the fully withdrawn position. Toward the end of phase three, the reactor sees a rapid increase in power that ends at $15\text{\,}\mathrm{s}$ when all control rods are forced into the reactor, ending the accident. In the fourth and final phase, reactor power decays as the delayed neutron population dies out. Figure 10 on the bottom shows plots of average flux in the $x$-$y$ plane at various points in time, including at $14.95\text{\,}\mathrm{s}$, the maximum power excursion. These results are from a simulation with ${10}^{6}$ particles in $10$ batches (a total of ${10}^{7}$ particles).
@@
Figure 10: Modified C5G7 stuck rod accident simulation: (top) flux densities through time showing the four phases shaded as gray, green, red, and blue respectively; (bottom) scalar flux on $x$-$y$ plane (top view) at points in time.
@@
We first modeled this problem using the seven-group materials in the C5G7 benchmark description [hou2017oecd], which we call C5G7 in the remainder of this work. We use $3.9\text{\times}{10}^{6}$ mesh cells in a 3D, time- and energy-dependent tally mesh. The movement of control rods into and out of the reactor is modeled with the continuous movement functionality of MC/DC [variansyah_2023_highfidelity]. To verify the delta tracking methods we explore convergence to the correct solutions using continuously moving surfaces, and we compare delta tracking solutions to solutions provided by surface tracking. This further verifies that delta tracking can be used in conjunction with continuously moving surfaces. Performance data is collected using ${10}^{6}$ particles in $10$ batches (total of ${10}^{7}$ particles).
@@ 5.3 C5CE problem
Next, we define a continuous-energy version of the previous problem (named C5CE), in which the reactor undergoes the same four-phase accident. We use the same C5G7 geometry [c5g7_zenodo], source distribution, and initial condition but establish each material using continuous energy cross sections. Table LABEL:tab:c5ce in Appendix B includes the material compositions for the C5CE problem. Using this benchmark, we evaluate both the voxelized tally method and the hybrid-in-energy method described in Section 3.3. Figure 1 shows the macroscopic total and majorant cross sections as a function of energy for the materials in C5CE. The neutron resolved-resonance region ends around $10\text{\,}\mathrm{keV}$, so we set the transition energy at $50\text{\,}\mathrm{keV}$. Particles moving at energies above $50\text{\,}\mathrm{keV}$ will use delta tracking, and those below will use surface tracking.
@@
5.4 Modified Dragon-burst problem
Conducted in 1944 during the Manhattan Project, the Dragon-burst experiments [kimpland_2021_dragon] proved that criticality and supercriticality could be achieved with prompt neutrons alone. Previous experiments, such as the Chicago Pile One, depended on the presence of delayed neutrons to achieve criticality. Figure 9 on the right shows a schematic of the experiment where a highly enriched (75%) uranium hydride slug (UH10) was dropped through a tamper with additional fuel. The slug moves through the core, triggering a prompt supercritical reaction. The reactivity excursion ends when gravity causes the slug to fall out of the core. Kimpland et al. showed that this burst criticality experiment achieved an increase in the neutron population of more than nine orders of magnitude [kimpland_2021_dragon]. The total time of the reaction is fast enough such that cross section heating can be neglected [kimpland_2021_dragon]. Here we consider a less-reactive version of the Dragon burst problem (25% enrichment) to test delta tracking with the hybrid-in-material method described in Section 3.2, and to provide additional verification for delta tracking with continuously moving surfaces. Figure 11 shows the overall flux density as a function of time on the top and a $y$-$z$ plot of the scalar flux at various points in time on the bottom.
Figure 11: Modified Dragon burst simulation: (top) flux density as a function of time, (bottom) scalar flux on $y$-$z$ plane (side view) at points in time.
@@ 5.4 Modified Dragon-burst problem
We expect delta tracking algorithms to perform quite poorly in this model for a number of reasons. First, when undergoing delta tracking, particles in the air region will experience a large number of rejected virtual collisions, since the majorant is orders of magnitude larger than the cross section of air. Second, as much of the problem involves a near-void material, the collision estimator typically required by delta tracking will generate poor tallies in those regions. Third, the problem is geometrically simple, consisting of a rectangular slug moving through a rectangular slab with a properly sized hole such that the slug can move through the slab. The benefit of delta tracking over surface tracking is most pronounced in problems with many surface crossings [woodcock_1965_deltatracking]. Our objectives of comparing delta tracking methods for this specific problem are to confirm that delta tracking works with such a dynamic problem with continuously moving materials and to assess whether the hybrid-in-material method outperforms standard delta tracking.
@@ 6 Results
In this section, we discuss the performance results of the benchmarks modeled with surface, delta, and voxelized delta tracking, as well as the hybrid-in-energy and hybrid-in-material tracking algorithms. We also verify that the various delta tracking methods converge to the same solution as surface tracking when transporting on a geometry with continuously moving surfaces. All data, plots, the exact version of MC/DC, and input decks are available from Zenodo [delta_restuls].
@@ 6.1 Performance results
All results presented below are executed on the Dane and Lassen high-performance computing systems at Lawrence Livermore National Laboratory (LLNL). Dane is a CPU-only system with dual-socket Intel Xeon Sapphire Rapids CPUs, each with 56 cores for a total of 112 CPU cores per node. Lassen is the open collaboration sibling to the Sierra machine with four Nvidia Tesla V100 GPUs and two IBM Power 9 CPUs per node. We make all performance statements with respect to a whole node of Dane (112 MPI threads, CPU) and a whole node of Lassen (4 MPI threads, GPU). We compile to CPUs with Numba v0.60.0, and compile to Nvidia GPUs with CUDA v11.8 and Nvidia-PTX with Numba v0.59.0666Numba v0.59.1 is the most recent version to support Power9 CPUs.. The delta tracking methods are built using MC/DC v0.12.0 [transport_cement_mcdc_2024], compiling on GPUs with Harmonize v0.0.2 [harmonize]. All floating-point operations use double precision.
@@
The performance of the methods on the modified Dragon problem is at odds with observations of the algorithms on the other problems. Delta tracking simulations (both standard and voxelized) do not reach the end of the specified end time on Dane or Lassen in the eight hours allocated, on either machine. On Dane, as predicted, the hybrid-in-material method improves performance over the other two methods, as it completes the simulation in about four hours. However, standard surface tracking requires only one hour for this simulation. The hybrid-in-material method did not finish on Lassen. These observations are consistent with the understanding that delta tracking is typically less effective for problems with fewer surfaces and significant material heterogeneities.
@@ 7 Discussion, Conclusions, and Future Work
We have also demonstrated a novel hybrid-in-energy surface-delta tracking scheme, where surface and delta tracking are used at low and high energies, respectively. On the continuous energy version of the C5G7 benchmark geometry undergoing a four-phase transient, we observe an order of magnitude increase in FOM with the hybrid-in-energy scheme, on both CPU and GPU nodes. We have also confirmed that hybrid surface-delta tracking methods can yield improved results in problems with significant void regions [leppanen_2010_burnup].
@@
The main takeaway of this work is that delta and surface tracking do not have to be treated as discrete choices in the numerical method, as corroborated by papers from the developers of Serpent2 and MONK/MCBEND  [leppanen_2017_collision, richards_monk_2015]. Greater performance can potentially be achieved by mixing and matching the underlying transport methods, given the physical parameters of a simulation and the relative strengths and weaknesses of a given transport application. As discussed in Section 3.4, implementing delta tracking in a surface tracking code is relatively simple, given a method to calculate a macroscopic majorant cross section.
@@
MC/DC does not currently treat the unresolved resonance region. Work is ongoing to support unresolved resonances; when implemented the energy cutoff for the hybrid-in-energy delta tracking method could use the average value of the cross section to determine a cutoff value. For the majorant determination in the unresolved resonance region, the maximum cross section value in the unresolved resonance table will be used. This work only includes analysis for fixed source time-dependent problems. We leave the analysis of $k$ or $\alpha$ eigenvalue calculations for future work.
Methods for reducing the dimensionality of the majorant are also being explored to achieve a more-efficient majorant cross section lookup in distance to collision operations. Experiments with the collision estimator for scalar flux and methods of tallying into vacuum and low interaction rate regions are being considered, including implementing a cutoff method [leppanen_2010_burnup] or weighted delta tracking [morgan_weighted-delta-tracking_2015].
@@ Appendix A Continuous Energy Macroscopic Majorant
To compute a unified energy grid per material, we combine the energy grids from all nuclides in a given material. Then we call numpy.unique(), which returns a sorted array (from smallest to largest) with no repeating elements [van_der_walt_numpy_2011]. To compute a unified energy grid for the whole problem, we do the same but with all the nuclides in the entire problem.
Computing a macroscopic majorant for nuclides that are not on a unified energy grid requires two levels of interpolation to put a given macroscopic total cross section on a unified energy grid. First, interpolae from each nuclide’s microscopic cross section onto a material’s unified energy grid to compute a macroscopic total cross section. Then, perform a second interpolation from the material to the majorant’s unified energy grid. We use scipy.interpolate.interp1d() to interpolate from one energy grid to the next [2020SciPy-NMeth].